# ORBGRAND Is Exactly Capacity-achieving via Rank Companding

### v1

@@ C-C Proof of Lemma 9
$\frac{1}{N}\ln\mathbb{E}\Bigl\{e^{N\theta\mathsf{D}(w^{\prime})}\Big\lvert\underline{\mathsf{T}}\Bigr\}$ $=\frac{1}{N}\sum_{n=1}^{mN}\ln\Bigl(1+e^{\frac{\theta}{m}\cdot\frac{n}{mN}}\Bigr)-m\ln 2$
@@
Appendix D Equivalence Between BICM Decoder (23) and SGRAND
@@
$\quad\Bigl\lvert\ln\frac{q_{j}^{+}(y_{i})}{q_{j}^{-}(y_{i})}\Bigr\rvert\cdot\mathbf{1}\Bigl(\text{sgn}\Bigl(\ln\frac{q_{j}^{+}(y_{i})}{q_{j}^{-}(y_{i})}\Bigr)\cdot x_{i,j}(w)<0\Bigr).$ (48)
@@ E-A Proof of Theorem 2
$\lim\limits_{N\to\infty}\mathbb{E}\mathsf{D}(1)$ $=\frac{1}{m}\sum_{j=1}^{m}\Big(\frac{1}{2}\int_{\mathcal{R}_{1,j}}\Bigl\lvert\ln\frac{q_{j}^{+}(y)}{q_{j}^{-}(y)}\Bigr\rvert q_{j}^{+}(y)\mathrm{d}y$
$\quad+\frac{1}{2}\int_{\mathcal{R}_{2,j}}\Bigl\lvert\ln\frac{q_{j}^{+}(y)}{q_{j}^{-}(y)}\Bigr\rvert q_{j}^{-}(y)\mathrm{d}y\Big).$ (54)
@@
$\lim\limits_{N\to\infty}\frac{1}{N}\ln\mathbb{E}\left\{e^{N\theta\mathsf{D}(w^{\prime})}\Big\lvert\underline{\mathsf{T}}\right\}$
$=\sum_{j=1}^{m}\mathbb{E}\Bigl[\ln\Bigl(1+e^{\frac{\theta}{m}\cdot\Bigl\lvert\ln\frac{q_{j}^{+}(\mathsf{Y})}{q_{j}^{-}(\mathsf{Y})}\Bigr\rvert}\Bigr)\Bigr]-m\ln 2.$ (56)
@@ E-B Proof of Theorem 3
$\lim\limits_{N\to\infty}\mathbb{E}\mathsf{D}(1)$ $=\frac{1}{m}\sum_{j=1}^{m}\Big(\frac{1}{2}\int_{\mathcal{R}_{1,j}}\bar{\Psi}\Bigl(\Bigl\lvert\ln\frac{q_{j}^{+}(y)}{q_{j}^{-}(y)}\Bigr\rvert\Bigr)q_{j}^{+}(y)\mathrm{d}y$
$\quad+\frac{1}{2}\int_{\mathcal{R}_{2,j}}\bar{\Psi}\Bigl(\Bigl\lvert\ln\frac{q_{j}^{+}(y)}{q_{j}^{-}(y)}\Bigr\rvert\Bigr)q_{j}^{-}(y)\mathrm{d}y\Big).$ (61)
@@
$\lim\limits_{N\to\infty}\frac{1}{N}\ln\mathbb{E}\left\{e^{N\theta\mathsf{D}(w^{\prime})}\Big\lvert\underline{\mathsf{T}}\right\}$
@@
In contrast to ${I}_{\text{ORB,BICM}}$ given by (26), here the expression (64) suggests that coding is done separately over different bit channels, rather than across them jointly. Only when all bit channels have identical laws, i.e., $q_{j}^{\pm}(y)$ being the same for $j=1,\ldots,m$, ${I}_{\text{ORB,BICM}}$ and $\tilde{I}_{\text{ORB,BICM}}$ coincide. Nevertheless, as shown in our numerical results in Section IV-E, their discrepancy is typically negligible.
@@ Appendix F Supporting Lemmas
$-\Bigl(\frac{1}{2}\int_{\mathcal{R}_{1}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{+}(y)\mathrm{d}y+\frac{1}{2}\int_{\mathcal{R}_{2}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{-}(y)\mathrm{d}y\Bigr)^{2}$ (65)

### v2

@@ C-C Proof of Lemma 9
$\frac{1}{N}\ln\mathbb{E}\Bigl\{e^{N\theta\mathsf{D}(w^{\prime})}\Big|\underline{\mathsf{T}}\Bigr\}$ $=\frac{1}{N}\sum_{n=1}^{mN}\ln\left(1+e^{\frac{\theta}{m}\cdot\frac{n}{mN}}\right)-m\ln 2$
@@
Appendix D Equivalence of BICM Decoder (23) and SGRAND
@@
$\quad\Bigl\lvert\ln\frac{q_{j}^{+}(y_{i})}{q_{j}^{-}(y_{i})}\Bigr\rvert\cdot\mathbf{1}\left(\text{sgn}\left(\ln\frac{q_{j}^{+}(y_{i})}{q_{j}^{-}(y_{i})}\right)\cdot x_{i,j}(w)<0\right).$ (48)
@@ E-A Proof of Theorem 2
$\lim\limits_{N\to\infty}\mathbb{E}\mathsf{D}(1)$ $=\frac{1}{m}\sum_{j=1}^{m}\Biggl(\frac{1}{2}\int_{\mathcal{R}_{1,j}}\Bigl\lvert\ln\frac{q_{j}^{+}(y)}{q_{j}^{-}(y)}\Bigr\rvert q_{j}^{+}(y)\mathrm{d}y$
$\quad+\frac{1}{2}\int_{\mathcal{R}_{2,j}}\Bigl\lvert\ln\frac{q_{j}^{+}(y)}{q_{j}^{-}(y)}\Bigr\rvert q_{j}^{-}(y)\mathrm{d}y\Biggr).$ (54)
@@
$\lim\limits_{N\to\infty}\frac{1}{N}\ln\mathbb{E}\left\{e^{N\theta\mathsf{D}(w^{\prime})}\Big|\underline{\mathsf{T}}\right\}$
$=\sum_{j=1}^{m}\mathbb{E}\Biggl[\ln\Biggl(1+e^{\frac{\theta}{m}\cdot\bigl\lvert\ln\frac{q_{j}^{+}(\mathsf{Y})}{q_{j}^{-}(\mathsf{Y})}\bigr\rvert}\Biggr)\Biggr]-m\ln 2.$ (56)
@@ E-B Proof of Theorem 3
$\lim\limits_{N\to\infty}\mathbb{E}\mathsf{D}(1)$ $=\frac{1}{m}\sum_{j=1}^{m}\Bigl(\frac{1}{2}\int_{\mathcal{R}_{1,j}}\bar{\Psi}\Bigl(\Bigl\lvert\ln\frac{q_{j}^{+}(y)}{q_{j}^{-}(y)}\Bigr\rvert\Bigr)q_{j}^{+}(y)\mathrm{d}y$
$\quad+\frac{1}{2}\int_{\mathcal{R}_{2,j}}\bar{\Psi}\Bigl(\Bigl\lvert\ln\frac{q_{j}^{+}(y)}{q_{j}^{-}(y)}\Bigr\rvert\Bigr)q_{j}^{-}(y)\mathrm{d}y\Bigr).$ (61)
@@
$\lim\limits_{N\to\infty}\frac{1}{N}\ln\mathbb{E}\left\{e^{N\theta\mathsf{D}(w^{\prime})}\Big|\underline{\mathsf{T}}\right\}$
@@
In contrast to ${I}_{\text{ORB,BICM}}$ given by (26), here the expression (64) suggests that coding is done separately over different bit channels, rather than across them jointly. When all bit channels have identical laws, i.e., $q_{j}^{\pm}(y)$ being the same for $j=1,\ldots,m$, ${I}_{\text{ORB,BICM}}$ and $\tilde{I}_{\text{ORB,BICM}}$ coincide. Nevertheless, as shown in our numerical results in Section IV-E, their discrepancy is typically negligible.
@@ Appendix F Supporting Lemmas
$-\left(\frac{1}{2}\int_{\mathcal{R}_{1}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{+}(y)\mathrm{d}y+\frac{1}{2}\int_{\mathcal{R}_{2}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{-}(y)\mathrm{d}y\right)^{2}$ (65)